OpenStudy (beautifulmystery.):

which decimal is equilvent to 5 9/20 @fatman96

OpenStudy (anonymous):

5.20

OpenStudy (beautifulmystery.):

thnx

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

your preety

OpenStudy (anonymous):

Don't listen to sameri, they're a troll

OpenStudy (anonymous):

your just mad because i got that question right

OpenStudy (anonymous):

This one? No you got this one wrong too.

OpenStudy (anonymous):

oh ok smart one

OpenStudy (anonymous):

Nah I'm not really smart, you're just a troll lol

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

keep saying stuff to make you feel better

OpenStudy (beautifulmystery.):

please leave this post

OpenStudy (anonymous):

\[\large \sf 5 \times \frac{9}{20}= \frac{45}{100}= .45\]

OpenStudy (beautifulmystery.):

thank you @LegendarySadist

OpenStudy (beautifulmystery.):

this really wasn't for me though

OpenStudy (anonymous):

No problem. So the answer is 5.45

OpenStudy (beautifulmystery.):

thnx!